RHIC Running: Nonlinear chromaticity corrections yesterday. Not much improvement.
Running at high per bunch intensities, 1.8e11 with polarizations of 50/45 at store. 55/50 at injection. PHENIX ZDC rates were as high as 14.5kHz. Also, in the middle of the night, the PHENIX ZDC luminosity increased by 15% for no apparent reason. STAR ZDC may have seen a slight increase at the same time, but not in the BBC.
There was a cryo lead flow fault that kept RHIC power supplies interlocked off for several hours.
It was discovered that the D1 extraction bump used for NSRL disrupts the polarized proton delivery to the AGS. It was turned off without any apparent effect on NSRL.
Today there will be APEX studies all day, 9:00 - 16:00. The end of APEX will be devoted to looking at the effect of cryo pumping on the 10Hz orbit oscillations. The cryo will be turned off at 3:00PM, and after ~45 minutes to let the system equilibrate, a measurement of the orbit will be taken with BPMs to see the strength of the 10Hz oscillations. When the measurements are completed, the cryo will be turned back on. It is hoped this measurement will provide input to the new working point problem.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |